# A differentiated replication study on LLMs

### v1

@@ Abstract
Index Terms:
Automated Vulnerability Repair, security
@@ I Introduction
Large language models (LLMs) are increasingly proposed for automated vulnerability repairs (AVR) [1]. At the time of submission, Google Scholar reports 483 works on the topic from 2024, excluding ArXiv papers, SoKs and systematic literature reviews. LLMs results (e.g., [2]) seem more promising than traditional repair approaches based on testing (e.g., [3]). APR4Vul tested traditional program repair methods on the Vul4J dataset [4] and obtained at best 5 correct patches. LLM methods obtained much better results: Codex was able to fix 10 vulnerabilities without fine-tuning [5], later works provided 12 perfect fixes [6], and up to 14 patches [7].
Yet, replicating LLMs’ success is challenging [1], mirroring issues previously observed with deep learning models [8]. We argue that three “invisible hands” could be at play: code leakage, completion vs patching, and perfect localization. These factors could help LLMs achieve “exceptional” results that testing-based program repair methods did not achieve.
The first “hand” involves code leakage: as a benchmark for vulnerability repair is eventually published, the new LLMs are likely to be trained on that very dataset, and the LLMs used for the original study might no longer be available. For instance, replicating the studies in [5, 9] require access to Codex, which is now deprecated. Further, suppose we wanted to use PrimeVul [10], a high-quality dataset with robust, albeit automatically verified, vulnerable and fixed functions. The corresponding paper was published at ICSE in May 2025, and therefore, GPT-4o-mini, deployed in May 2024, would seem to be a good candidate (it is used in 34 of the mentioned LLM papers). Unfortunately, the dataset was published on GitHub in March 2024 and includes commits pushed even earlier, so the possibility of data leakage cannot be ruled out.
The second and third invisible hands stem from perfect fault localization and prompt design. Providing the exact location of the vulnerability is often the gold standard in program repair, e.g., in [3] (non-LLM) and [11] (LLM-based). By itself, this is a reasonable approach to distinguish success in vulnerability localization from success in patch generation. The key problem happens when it is combined with prompt designs that frame repairs as code completion: the fragment to be patched is replaced with a placeholder, and the LLM is asked to complete the code (e.g. [9, 7, 5]). Given the likely leakage of code (see first invisible hand), LLMs may simply reproduce the correct fix via completion, a task where LLMs excel [12].
@@
Our is a skeptical replication study, aimed at disentangling the potential influence of “invisible hands” on the success of LLM-based AVR. How can we remove the influence of these hands in order to test the hypothesis? We need to measure some events that are very unlikely to happen if the LLM is generalizing (“ideal” LLM) vs the LLM is just relying on memorization (“cheater” LLM). Bayes theorem would then allows us to distinguish the posterior probability. Some preliminary evidence in this respect has been already explored for normal bug fixing [13]. Our idea is to systematically introduce errors in the localization of vulnerabilities.
@@
To strengthen the result, we add a second step: we ask a second LLM to review the result of the first LLM before testing. This is considered a good solution to scale [14] albeit a recent theoretical analysis [15] has shown that MLs patch-reviewer model can only be useful if precision and recall are above a certain threshold. Worse, if our overarching hypothsis is true the second LL M is likely to suffer from the same memorization issues that the first LLM is being tested. If LLMs merely reproduce and recognize fixes they have already seen, we expect cheater LLMs as reviewers to recognize (memorized) plausible patches better than other plausible but wrong patches. As soon as we perturb the pipeline so that recognition is hampered, such difference will disappear. Localization rrors will not make a difference as they are ignored by cheaters.
Operationally, we build on the pipelines from [5] and [7]: we assess the impact of different prompt designs and include in our input vulnerabilities obtained through code transformations so that they have no corresponding published fix. This contribute to isolate model generalization from memorization. We will validate generated patches using both regression tests and Proof of Vulnerability (PoV) tests [^1] A proof of vulnerability in test-based vulnerability repair [16, 3, 17] is a test which flips between the vulnerable and non-vulnerable code fragments.
@@ II Terminology and background
Automated Program Repair (APR) aims to support developers by automatically fixing bugs in software. End-to-end APR pipelines cover the full repair process: locating the faulty code (fault localization), generating a patch, and validating it to ensure correctness [18]. The ultimate goal is to produce patches that can be reviewed and accepted by developers.
To evaluate APR tools, researchers use benchmarks composed of buggy programs and associated test suites. A patch is considered plausible if it passes all available tests automatically. However, only correct patches are semantically equivalent to the original developer fix, while overfitted patches pass tests but fail to fix the root cause of the bug [16, 3, 5].
Automated Vulnerability Repair (AVR) is a specialization of APR focused on fixing security-related bugs—vulnerabilities that attackers could exploit. In addition to preserving functionality, AVR patches must restore or maintain security properties. As a result, AVR benchmarks often include not only unit tests but also Proof of Vulnerability (PoV) tests [4], which demonstrate that the original bug can be exploited.
In the literature, AVR-generated patches that pass all tests are also referred to as plausible, as in general APR [5], or as End-to-End (E2E) tested patches [3]. Upon manual inspection, such patches may turn out to be correct, overfitted, or security-fixing—patches that remove the vulnerability but silently introduce regressions that break functionality [3].
@@ III-1 Evaluation pipelines for AVR
Evaluating tools for repairing vulnerabilities required the development of benchmarks provided with functional and security tests. Pinconschi et al. [16] developed SecureThemAll to test APR techniques for repairing 55 C/C++ security faults from the DARPA Challenge Sets [19], which approximate real-world vulnerabilities and include functional and PoV tests. The 10 APR tools tested received only C source code and its test suite. Bui et al. [3] proposed APR4Vul to evaluate general APR techniques on vulnerabilities in Vul4J [4], a manually curated benchmark of 79 reproducible Java vulnerabilities, each with functional and PoV tests. APR4Vul tools were provided with exact, manually defined vulnerability locations to isolate the patch generation task from localization errors.
@@ III-2 Evaluations of LLMs for AVR
Different studies proposed testing LLM to repair software vulnerabilities. Pearce et al. [9] tested proprietary and open-source LLMs, including a locally trained model, for zero-shot vulnerability repair in C/C++. They evaluated these against synthetic vulnerabilities and 12 real-world CVEs using six prompt templates (Table I) to present the exact vulnerability location. Notably, prompts like “n.h.”, “s.1”, and “s.2” omit the faulty code, framing the task as code generation rather than repair—contrary to AVR’s assumption that the defect is known.
TABLE I: Prompt templates used by Pearce et al. [9].
@@
Wu et al. [5] proposed a framework to test DL-based tools for AVR in Java. They extended the Vul4J benchmark [4] and proposed VJBench-trans, a new dataset of transformed versions of the collected single-hunk vulnerabilities. The transformation should mitigate the advantage for LLMs that were already exposed to the testing data during their training. The prompt templates that the authors use are reported in Table II, and all of them but “Codex” present a code completion rather than a vulnerability repair problem.
TABLE II: Prompt templates used by Wu et al. [5].
@@
CodeT5 Mask buggy lines with $<$extra_id_0$>$ and input the buggy function
@@
PLBART Mask buggy lines with $<$mask$>$ and input the buggy function
InCoder Mask buggy lines with $<$mask$>$ and input the buggy function
@@
Figure 1: While different studies tested LLMs for repair starting from the exact vulnerability localization, we investigate the impact of errors in the localization for different approaches. For example, in the approach proposed by Wu et al. [5], the LLM is prompted to substitute exactly the vulnerable lines, so even small displacements would prove disruptive. Here we prompted GPT3.5 (used by Kulsum et al. [7]) to fix the vulnerable function, but gave in the prompt wrong information about the vulnerable line. In the figure, we have two negative and one positive offsets and the corresponding responses. When the offset is 4 lines above, pointing to just a curly bracket (line 2), the model still generates the developer fix. How is this likely to happen in the absence of memorization?
Kulsum et al. [7] replicated the work by Pearce et al. with GPT‑3.5 [20] and added chain-of-thought prompts and iterative feedback by external tools to improve the repair process. They tested their approach against 60 real-world CVEs (10 in C and 50 in Java from the Vul4J [4] and VJBench [5] benchmarks). Their first prompt follow the same formats proposed by Pearce et al.
Fu et al. [21] proposed AIBugHunter, an ML-based software vulnerability analysis tool for C/C++ for Visual Studio Code. They integrate their tools LineVul [22] for vulnerability localization and VulRepair [23] for vulnerability repair. However, the pipeline is only evaluated via a qualitative user study, making its objective performance hard to assess.
@@ III-3 LLM second opinion
As code reviewing is a tedious and expensive process for software development, different studies have proposed leveraging LLMs to ease it. Jensen et al. [24] investigated using LLMs in reviews to evaluate both code security and functionality. The performance of proprietary models in particular is impressive (over 95% accuracy for vulnerability detection, over 88% F1 score for functionality validation), however it is not clear to what extent this could be due to data leakage. Recent works from Jaoua et al. [25] and Kavian et al. [26] combine the use of static analyzers and LLMs to review and improve the quality of developer and LLM-generated patches, respectively.
@@ IV Research Statement
Figure 2 summarizes our research questions, which start from the differentiated replication of the setups in “How Effective Are Neural Networks for Fixing Security Vulnerabilities” [5] and VRPilot [7]. Building on the mentioned studies, we use different prompt templates to investigate how they affect the repair capabilities of LLMs.
  1. $\mathbf{RQ1}$ Reproducibility baseline. How many vulnerabilities can LLMs fix when provided with their exact localization and how sensitive to the prompt are the results?
@@
To scientifically validate this hypothesis we cannot use the traditional approach used in software engineering which tests for significant differences. We claim that two treatments (showing the line or providing vulnerability information) have the same effect. This is the same idea behind statistically proving that a generic label drug has the same effect of a branded label drug [27]. We need to use a different approach, and namely prove significant equivalence [28] and not just no significant difference. In this set-up our alternative hypothesis are hypothesis about equivalence.
@@
Prompting an LLM with an incorrect vulnerability localization may lead to two possible outcomes: the model might ignore the incorrect line and regenerate the known fix, or it might fail to recognize the original fixing pattern and produce an incorrect patch — even in cases where it previously succeeded with accurate localization. If the second outcome plays a significant role, we would expect to see a decline in performance when localization errors are introduced. We hypothesize that any deviation from the correct line will be enough to trigger this effect, and therefore we do not expect meaningful performance differences between small (2$\leq$ 2≤ 2) and larger (444, 888) displacements. From 0 and 1 the difference will not be enough because we cannot distinguish the “honest” and the “cheating” because of the git diff window.
$H^{alt}_{\mbox{equiv-err-order}}$: There is statistically significant equivalence in the performance between patches generated with prompts for which the line-level location is shifted by a small (<3<3< 3 in magnitude) and a considerable (>3>3> 3) offset.
@@
  1. $\mathbf{RQ2}$ Second opinion. How many incorrect patches can LLMs detect before the testing phase?
@@
  1. $\mathbf{RQ2_{ERR}}$ Second opinion with error. How many incorrect patches can LLMs detect before the testing phase?
@@
  1. $\mathbf{RQ3}$ Manual inspection. How many patches that pass tests are semantically correct?
We will also investigate whether the justification of the LLM is also correct, following the preliminary findings of [29].
@@ V-1 Dataset
  • D2: The dataset contains real-world data, as synthetic data could distort performances (see the difference on synthetic and real-world data by Pearce et al. [9]).
@@
  • D5: The dataset contains refactored vulnerabilities (e.g.,control flow has been changed [5] or identifiers are renamed in a language that is not English), so that the corresponding developer fix is not available.
For the listed criteria, our choice falls on the dataset VJBench and VJBench-trans [5] built to extend the benchmark Vul4J [4]. These datasets contain Java vulnerabilities, of which at least 50 single-hunk vulnerabilities (Wu et al. selected them in 2023, but Vul4J was expanded since then). We aim to widen this selection to single-function vulnerabilities and use the scripts of the refactoring process of VJBench-trans to generate vulnerabilities that the LLMs have never seen.
@@ V-2 Prompts
Table III summarizes the four prompt templates we plan to use, building on previous works that evaluated LLMs for AVR [9, 5] with two levels of information:
@@
In all cases we provide the code of the vulnerable function, but in the first one we add no additional information (as if the vulnerability localization was performed by ReVeal [30]), the second provides line-level localization (like LineVul [22] could do), the third one provides the vulnerability type (as ChatGPT was prompted to do in [31]) and the last one both line-level localization and vulnerability type (as a SAST tool could do [32]). Besides the code and eventually the vulnerability type, each prompt will include a request to generate a new version of the given function to fix the present vulnerability. If other prompts are equivalent to the bare one, we know they all memorize and the additional information is irrelevant.

### v2

@@ I Introduction
Large language models (LLMs) are increasingly proposed for automated vulnerability repairs (AVR) [zhou2024large]. At the time of submission, Google Scholar reports 483 works on the topic from 2024, excluding ArXiv papers, SoKs and systematic literature reviews. LLMs results (e.g., [huang2025comprehensive]) seem more promising than traditional repair approaches based on testing (e.g., [bui2024apr4vul]). APR4Vul tested traditional program repair methods on the Vul4J dataset [bui2022vul4j] and obtained at best 5 correct patches. LLM methods obtained much better results: Codex was able to fix 10 vulnerabilities without fine-tuning [wu2023effective], later works provided 12 perfect fixes [sagodi2024reality], and up to 14 patches [kulsum2024case].
Yet, replicating LLMs’ success is challenging [zhou2024large], mirroring issues previously observed with deep learning models [sejfia2024toward]. We argue that three “invisible hands” could be at play: code leakage, completion vs patching, and perfect localization. These factors could help LLMs achieve “exceptional” results that testing-based program repair methods did not achieve.
The first “hand” involves code leakage: as a benchmark for vulnerability repair is eventually published, the new LLMs are likely to be trained on that very dataset, and the LLMs used for the original study might no longer be available. For instance, replicating the studies in [wu2023effective, pearce2023examining] require access to Codex, which is now deprecated. Further, suppose we wanted to use PrimeVul [ding2024vulnerability], a high-quality dataset with robust, albeit automatically verified, vulnerable and fixed functions. The corresponding paper was published at ICSE in May 2025, and therefore, GPT-4o-mini, deployed in May 2024, would seem to be a good candidate (it is used in 34 of the mentioned LLM papers). Unfortunately, the dataset was published on GitHub in March 2024 and includes commits pushed even earlier, so the possibility of data leakage cannot be ruled out.
The second and third invisible hands stem from perfect fault localization and prompt design. Providing the exact location of the vulnerability is often the gold standard in program repair, e.g., in [bui2024apr4vul] (non-LLM) and [li2024hybrid] (LLM-based). By itself, this is a reasonable approach to distinguish success in vulnerability localization from success in patch generation. The key problem happens when it is combined with prompt designs that frame repairs as code completion: the fragment to be patched is replaced with a placeholder, and the LLM is asked to complete the code (e.g. [pearce2023examining, kulsum2024case, wu2023effective]). Given the likely leakage of code (see first invisible hand), LLMs may simply reproduce the correct fix via completion, a task where LLMs excel [zhang2024instruct].
@@
Our is a skeptical replication study, aimed at disentangling the potential influence of “invisible hands” on the success of LLM-based AVR. How can we remove the influence of these hands in order to test the hypothesis? We need to measure some events that are very unlikely to happen if the LLM is generalizing (“ideal” LLM) vs the LLM is just relying on memorization (“cheater” LLM). Bayes theorem would then allows us to distinguish the posterior probability. Some preliminary evidence in this respect has been already explored for normal bug fixing [kong2025demystifying]. Our idea is to systematically introduce errors in the localization of vulnerabilities.
@@
To strengthen the result, we add a second step: we ask a second LLM to review the result of the first LLM before testing. This is considered a good solution to scale [hou2024largesurvey] albeit a recent theoretical analysis [camporese2025using] has shown that MLs patch-reviewer model can only be useful if precision and recall are above a certain threshold. Worse, if our overarching hypothsis is true the second LL M is likely to suffer from the same memorization issues that the first LLM is being tested. If LLMs merely reproduce and recognize fixes they have already seen, we expect cheater LLMs as reviewers to recognize (memorized) plausible patches better than other plausible but wrong patches. As soon as we perturb the pipeline so that recognition is hampered, such difference will disappear. Localization rrors will not make a difference as they are ignored by cheaters.
Operationally, we build on the pipelines from [wu2023effective] and [kulsum2024case]: we assess the impact of different prompt designs and include in our input vulnerabilities obtained through code transformations so that they have no corresponding published fix. This contribute to isolate model generalization from memorization. We will validate generated patches using both regression tests and Proof of Vulnerability (PoV) tests [^1] A proof of vulnerability in test-based vulnerability repair [pinconschi2021comparative, bui2024apr4vul, zhang2022program] is a test which flips between the vulnerable and non-vulnerable code fragments.
@@ II Terminology and background
Automated Program Repair (APR) aims to support developers by automatically fixing bugs in software. End-to-end APR pipelines cover the full repair process: locating the faulty code (fault localization), generating a patch, and validating it to ensure correctness [shen2020survey]. The ultimate goal is to produce patches that can be reviewed and accepted by developers.
To evaluate APR tools, researchers use benchmarks composed of buggy programs and associated test suites. A patch is considered plausible if it passes all available tests automatically. However, only correct patches are semantically equivalent to the original developer fix, while overfitted patches pass tests but fail to fix the root cause of the bug [pinconschi2021comparative, bui2024apr4vul, wu2023effective].
Automated Vulnerability Repair (AVR) is a specialization of APR focused on fixing security-related bugs—vulnerabilities that attackers could exploit. In addition to preserving functionality, AVR patches must restore or maintain security properties. As a result, AVR benchmarks often include not only unit tests but also Proof of Vulnerability (PoV) tests [bui2022vul4j], which demonstrate that the original bug can be exploited.
In the literature, AVR-generated patches that pass all tests are also referred to as plausible, as in general APR [wu2023effective], or as End-to-End (E2E) tested patches [bui2024apr4vul]. Upon manual inspection, such patches may turn out to be correct, overfitted, or security-fixing—patches that remove the vulnerability but silently introduce regressions that break functionality [bui2024apr4vul].
@@ III-1 Evaluation pipelines for AVR
Evaluating tools for repairing vulnerabilities required the development of benchmarks provided with functional and security tests. Pinconschi et al. [pinconschi2021comparative] developed SecureThemAll to test APR techniques for repairing 55 C/C++ security faults from the DARPA Challenge Sets [CGC], which approximate real-world vulnerabilities and include functional and PoV tests. The 10 APR tools tested received only C source code and its test suite. Bui et al. [bui2024apr4vul] proposed APR4Vul to evaluate general APR techniques on vulnerabilities in Vul4J [bui2022vul4j], a manually curated benchmark of 79 reproducible Java vulnerabilities, each with functional and PoV tests. APR4Vul tools were provided with exact, manually defined vulnerability locations to isolate the patch generation task from localization errors.
@@ III-2 Evaluations of LLMs for AVR
Different studies proposed testing LLM to repair software vulnerabilities. Pearce et al. [pearce2023examining] tested proprietary and open-source LLMs, including a locally trained model, for zero-shot vulnerability repair in C/C++. They evaluated these against synthetic vulnerabilities and 12 real-world CVEs using six prompt templates (Table I) to present the exact vulnerability location. Notably, prompts like “n.h.”, “s.1”, and “s.2” omit the faulty code, framing the task as code generation rather than repair—contrary to AVR’s assumption that the defect is known.
TABLE I: Prompt templates used by Pearce et al. [pearce2023examining].
@@
Wu et al. [wu2023effective] proposed a framework to test DL-based tools for AVR in Java. They extended the Vul4J benchmark [bui2022vul4j] and proposed VJBench-trans, a new dataset of transformed versions of the collected single-hunk vulnerabilities. The transformation should mitigate the advantage for LLMs that were already exposed to the testing data during their training. The prompt templates that the authors use are reported in Table II, and all of them but “Codex” present a code completion rather than a vulnerability repair problem.
TABLE II: Prompt templates used by Wu et al. [wu2023effective].
@@
CodeT5 Mask buggy lines with <<extra_id_0>> and input the buggy function
@@
PLBART Mask buggy lines with <<mask>> and input the buggy function
InCoder Mask buggy lines with <<mask>> and input the buggy function
@@
Figure 1: While different studies tested LLMs for repair starting from the exact vulnerability localization, we investigate the impact of errors in the localization for different approaches. For example, in the approach proposed by Wu et al. [wu2023effective], the LLM is prompted to substitute exactly the vulnerable lines, so even small displacements would prove disruptive. Here we prompted GPT3.5 (used by Kulsum et al. [kulsum2024case]) to fix the vulnerable function, but gave in the prompt wrong information about the vulnerable line. In the figure, we have two negative and one positive offsets and the corresponding responses. When the offset is 4 lines above, pointing to just a curly bracket (line 2), the model still generates the developer fix. How is this likely to happen in the absence of memorization?
Kulsum et al. [kulsum2024case] replicated the work by Pearce et al. with GPT‑3.5 [openai2023gpt35] and added chain-of-thought prompts and iterative feedback by external tools to improve the repair process. They tested their approach against 60 real-world CVEs (10 in C and 50 in Java from the Vul4J [bui2022vul4j] and VJBench [wu2023effective] benchmarks). Their first prompt follow the same formats proposed by Pearce et al.
Fu et al. [fu2024aibughunter] proposed AIBugHunter, an ML-based software vulnerability analysis tool for C/C++ for Visual Studio Code. They integrate their tools LineVul [fu2022linevul] for vulnerability localization and VulRepair [fu2022vulrepair] for vulnerability repair. However, the pipeline is only evaluated via a qualitative user study, making its objective performance hard to assess.
@@ III-3 LLM second opinion
As code reviewing is a tedious and expensive process for software development, different studies have proposed leveraging LLMs to ease it. Jensen et al. [jensen2024software] investigated using LLMs in reviews to evaluate both code security and functionality. The performance of proprietary models in particular is impressive (over 95% accuracy for vulnerability detection, over 88% F1 score for functionality validation), however it is not clear to what extent this could be due to data leakage. Recent works from Jaoua et al. [jaoua2025combining] and Kavian et al. [kavian2024llm] combine the use of static analyzers and LLMs to review and improve the quality of developer and LLM-generated patches, respectively.
@@ IV Research Statement
Figure 2 summarizes our research questions, which start from the differentiated replication of the setups in “How Effective Are Neural Networks for Fixing Security Vulnerabilities” [wu2023effective] and VRPilot [kulsum2024case]. Building on the mentioned studies, we use different prompt templates to investigate how they affect the repair capabilities of LLMs.
  1. $\mathbf{RQ1}$ - Reproducibility baseline. How many vulnerabilities can LLMs fix when provided with their exact localization and how sensitive to the prompt are the results?
@@
To scientifically validate this hypothesis we cannot use the traditional approach used in software engineering which tests for significant differences. We claim that two treatments (showing the line or providing vulnerability information) have the same effect. This is the same idea behind statistically proving that a generic label drug has the same effect of a branded label drug [food2001guidance]. We need to use a different approach, and namely prove significant equivalence [meyners2012equivalence] and not just no significant difference. In this set-up our alternative hypothesis are hypothesis about equivalence.
@@
Prompting an LLM with an incorrect vulnerability localization may lead to two possible outcomes: the model might ignore the incorrect line and regenerate the known fix, or it might fail to recognize the original fixing pattern and produce an incorrect patch — even in cases where it previously succeeded with accurate localization. If the second outcome plays a significant role, we would expect to see a decline in performance when localization errors are introduced. We hypothesize that any deviation from the correct line will be enough to trigger this effect, and therefore we do not expect meaningful performance differences between small ($\leq 2$) and larger ($4$, 88) displacements. From 0 and 1 the difference will not be enough because we cannot distinguish the “honest” and the “cheating” because of the git diff window.
$H^{alt}_{\mbox{equiv-err-order}}$: There is statistically significant equivalence in the performance between patches generated with prompts for which the line-level location is shifted by a small (<3<3 in magnitude) and a considerable (>3>3) offset.
@@
  1. $\mathbf{RQ2}$ - Second opinion. How many incorrect patches can LLMs detect before the testing phase?
@@
  1. $\mathbf{RQ2_{ERR}}$ - Second opinion with error. How many incorrect patches can LLMs detect before the testing phase?
@@
  1. $\mathbf{RQ3}$ - Manual inspection. How many patches that pass tests are semantically correct?
We will also investigate whether the justification of the LLM is also correct, following the preliminary findings of [risse2025top].
@@ V-1 Dataset
D2: The dataset contains real-world data, as synthetic data could distort performances (see the difference on synthetic and real-world data by Pearce et al. [pearce2023examining]).
@@
D5: The dataset contains refactored vulnerabilities (e.g.,control flow has been changed [wu2023effective] or identifiers are renamed in a language that is not English), so that the corresponding developer fix is not available. For the listed criteria, our choice falls on the dataset VJBench and VJBench-trans [wu2023effective] built to extend the benchmark Vul4J [bui2022vul4j]. These datasets contain Java vulnerabilities, of which at least 50 single-hunk vulnerabilities (Wu et al. selected them in 2023, but Vul4J was expanded since then). We aim to widen this selection to single-function vulnerabilities and use the scripts of the refactoring process of VJBench-trans to generate vulnerabilities that the LLMs have never seen.
@@ V-2 Prompts
Table III summarizes the four prompt templates we plan to use, building on previous works that evaluated LLMs for AVR [pearce2023examining, wu2023effective] with two levels of information:
@@
In all cases we provide the code of the vulnerable function, but in the first one we add no additional information (as if the vulnerability localization was performed by ReVeal [chakraborty2021deep]), the second provides line-level localization (like LineVul [fu2022linevul] could do), the third one provides the vulnerability type (as ChatGPT was prompted to do in [fu2023chatgpt]) and the last one both line-level localization and vulnerability type (as a SAST tool could do [jasz2022end]). Besides the code and eventually the vulnerability type, each prompt will include a request to generate a new version of the given function to fix the present vulnerability. If other prompts are equivalent to the bare one, we know they all memorize and the additional information is irrelevant.